OpenStudy (anonymous):

I have the answer can somone please check me f(x)=(500x)/(x^2+1) has a relative maximum at x=? and a relative minimum at x=?

OpenStudy (agent0smith):

What were your answers? you can check by seeing the graph to get a general idea of where they should be: https://www.google.com/search?q=(500x)%2F(x%5E2%2B1)&aq=f&oq=(500x)%2F(x%5E2%2B1)&aqs=chrome.0.57&sourceid=chrome&ie=UTF-8

OpenStudy (anonymous):

so then you got 0 for the maximum and minimum right

OpenStudy (agent0smith):

x will not be zero.

OpenStudy (anonymous):

that is what I got so that is why I am questioning yall

OpenStudy (anonymous):

\[f'(x)={500\over x^2+1}-{1000x^2\over (x^2+1)^2}=0\\ 500(x^2+1)-1000x^2=0\\ 1-x^2=0\\ \boxed{x=1}\qquad\boxed{x=-1} \]

OpenStudy (anonymous):

one of them is the maxima, the other a minima.

OpenStudy (anonymous):

oh ok I understand what I did wrong so then the maximum is 1 mininmum is -1

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok thanks